# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

### v2

@@ 4 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
(4.1) $\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\left(\epsilon\lambda^{1/2}+5n^{2}\pi\log_{2}\frac{10r_{in}}{\epsilon}\right).$
@@
(4.2) $\leq C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}|.$
@@
and therefore $f(\Lambda)$ is monotone decreasing on $[\frac{1}{\mathrm{width}(\Omega)^{2}},\infty)$.
@@
For open sets, we also have a uniform estimate. Let us remark that, since the boundary of an open set may be complicate, in general one can not have a uniform estimate in term of its boundary as in Theorem 1.5; see [9, Theorem 3.14] for similar but earlier result.
Note that $C_{1}(n-1)$ is defined via (1.11).
Theorem 4.2.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded open domain, $n\geq 2$, let $Q$ be a cube that contains $\Omega$. For any $20<\epsilon<\ell(Q)/$ 2, it holds for all $k\in{\mathbb{N}}$ that
(4.3) $k\leq\frac{|\Omega_{\epsilon}|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}+\frac{\omega(n)\lambda^{\frac{n-1}{2}}}{(2\pi)^{n}}\left(2n^{2}\pi\frac{|Q\setminus\Omega_{\epsilon}|}{\epsilon}-C_{1}(n-1)\ell(Q)^{n-1}\right).$
where $\Omega_{\epsilon}=\{x\in Q:\,\mathrm{dist}(x,\Omega)<\epsilon\}$. In particular, one can take $Q$ to be of side length $\mathrm{diam}(\Omega)$.
Proof.
Since $\epsilon<\ell(Q)/2$, we can divide $\mathscr{R}$ into the union of cubes $Q_{j}$, which are cubes of side length $2^{-k}\ell(Q)$, where $k\in\mathbb{Z}_{+}$ is such that
(4.4) $\epsilon/2<2^{-k}\ell(Q)\sqrt{n}\leq\epsilon.$
Denote by $\tilde{Q}$ the union of all $Q_{j}$, such that $Q_{j}$ entirely locates in $Q\setminus\overline{\Omega}$. Then the number of $Q_{j}$ that composes $\tilde{Q}$ is controlled by
#$\#\{Q_{j}:\,Q_{j}\subset\tilde{Q}\}\leq\frac{|Q\setminus\Omega|}{2^{-kn}\ell(Q)^{n}}.$
On the other hand, $\Omega\subset\mathscr{R}\setminus\tilde{Q}$, which is further contained in
$\{x\in Q:\,\mathrm{dist}(x,\Omega)<2^{-k}\ell(Q)\sqrt{n}\}\subset\Omega_{\epsilon}.$
By Theorem 2.3, we have
$\mathcal{N}^{D}_{Q}(\lambda)$ $\leq\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)_{+}.$
This together with the monotonicity of Dirichlet eigenvalues implies that
$\frac{|Q|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\left(1-\frac{C_{1}(n-1)}{\ell(Q)\sqrt{\lambda}}\right)_{+}$ $\geq\mathcal{N}^{D}_{Q}(\lambda)\geq\mathcal{N}^{D}_{Q\setminus\tilde{Q}}(\lambda)+\mathcal{N}^{D}_{\tilde{Q}}(\lambda)$

### v3

@@ 4 Quantitative remainder estimate
$\mathcal{N}_{\Omega}^{D}(\lambda)$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda^{\frac{n-1}{2}}\left(\epsilon\lambda^{1/2}+5n^{2}\pi\log_{2}\frac{10r_{in}}{\epsilon}\right).$ (4.1)
@@
$\leq C_{Lip}(\Omega)\epsilon|\partial\Omega_{1}|.$ (4.2)
@@
and therefore $f(\Lambda)$ is monotone decreasing on $\frac{1}{\mathrm{width}(\Omega)^{2}},\infty)$.
@@
Acknowledgments
We wish to thank Prof. Rupert Frank and Prof. Simon Larson for bringing references [9, 12, 28] to us and several useful comments, and thank Dr. Xiang He for bringing the reference [32]. R. Jiang wishes to thank Prof. Xiao Zhong for explaining to him Pólya’s conjecture and helpful discussions at University of Jyväskylä around 2010. R. Jiang was partially supported by NNSF of China (12526205 & 12471094), F.H. Lin was in part supported by the National Science Foundation Grant DMS2247773 and DMS1955249.